Tuesday, July 13, 2010

Summary Night Report 20100713

Report for night 2010-07-13

SA: Encarni
SO: Patrick, Operonomer Hannah
Others: Deon, Hitesh, Morgan, Chris, Zolisa, Siphelo, Charl and the IQ team (Lisa, Francois, Darragh, James)

Notes and messages

* The last temp control unit for the aircon, which we installed last
night in the control room, is now dead. And it's freezing in here!!!! :(

* ELS issues:

During the day we had dew point warning on SOMMI which was clearly wrong.

Text messages are being sent by the ELS to the standby phone
which are irrelevant (SAMS comms lost, DEW point warnings
etc). This needs to be fixed before the standby person drowns the phone...

Night temp variance was yellow both nights, but the temps were
nicely equalized and temp gradients are flat as a pancake.

* Aircon for conditioning the building during the day does not work
due to low levels of glycol. Opening louvers early instead.

* Video feed from star tracker is working!! Thanks Charl and Hitesh!
Erm.. you do know I don't have a kingdom to swap for the camera,
right? ;)

* Tracker brakes release error three times. Twice it recovered
by itself, the other time it did not and did not move, although reporting
green and SLEW.

No comments:

Post a Comment